OpenStudy (anonymous):

if Drsuzuki rides his bike to his office, he averages 6 miles per hour. If he drives his car, he averages 12 miles per hour. His time driving is 1/2 hour less than his time bicycling.How far is his office?

OpenStudy (king):

1 mile

sam (.sam.):

Let t = time on the bike Then (t-0.5) = time in the car distance equation, Distance = Speed X time car distance = bike distance 12(t-0.5) =6t t = 1 hr Find distance using time on the bike: 1* 6 = 6 miles